# The Very High Energy Afterglow of Structured Jets: GW 170817 and Prospects for Future Detections

### v1

@@ 2.2 Dynamics of the deceleration
The structured outflow decelerates in an external medium, with an assumed density profile as a function of radius $R$:
@@
with $s=0$ for a constant-density external medium as considered in the following to model the afterglow of GW 170817. In this case we write $A=n_{\rm ext}m_{\rm p}$ with $n_{\rm ext}$ the external medium density. We focus on the dynamics of the shocked external medium at the forward shock. It is computed assuming that (i) the dynamics of each ring of material at angle $\theta$ is independent of other angles; (ii) the lateral expansion of the outflow is negligible. These two assumptions are questionable at late times, close to the transition to the Newtonian regime (Rhoads, 1997).
The Lorentz factor $\Gamma(\theta;R)$ of the shocked material at angle $\theta$ and radius $R$ is computed in a simplified way, using energy conservation (see e.g. Panaitescu & Kumar, 2000; Gill & Granot, 2018):
@@
with $M_{\mathrm{ext}}(R)=\int_{0}^{R}\rho_{\mathrm{ext}}(R)\,R^{2}\mathrm{d}R$ the swept-up mass per unit solid angle and $M_{\mathrm{ej}}(\theta)=\epsilon_{0}(\theta)/\Gamma_{0}(\theta)c^{2}$ the ejected mass per unit solid angle. The deceleration radius of the material at angle $\theta$ is defined by
@@
For a power-law structure as defined by Equations (58), the deceleration radius scales as $R_{\mathrm{dec}}(\theta)\propto\theta^{\frac{2b-a}{3-s}}$. Best-fit models of the afterglow of GW 170817 usually have $2b>a$ so that the deceleration radius is larger at high latitude (see § 4.2). More generally, Beniamini et al. (2020b) derived the conditions for the observed afterglow light curve to have a single peak, as in the case of GW 170817. For an off-axis observation with a large viewing angle, these conditions are $\Gamma_{0}^{\mathrm{c}}\theta_{\mathrm{c}}>1$; $2b>a/(4-s)$ (i.e. $8b>a$ for a uniform external medium); $b>-\ln{\Gamma_{0}^{\mathrm{c}}}/\ln{\theta_{\mathrm{c}}}$; and $\theta_{\mathrm{v}}>\theta_{\mathrm{c}}\left(\Gamma_{0}^{\mathrm{c}}\theta_{\mathrm{c}}\right)^{\frac{1}{b-1}}$.
This description allows us to characterize the early-time dynamics in the coasting phase $R\ll R_{\mathrm{dec}}(\theta)$ where the Lorentz factor remains constant. It branches continuously to the relativistic self-similar evolution (Blandford & McKee, 1976) for $R_{\mathrm{dec}}(\theta)\ll R\ll R_{\mathrm{N}}(\theta)=\Gamma_{0}^{\frac{2}{3-s}}(\theta)R_{\mathrm{dec}}(\theta)$, and finally to the Sedov-Taylor phase (Sedov, 1946; Taylor, 1950) in the non-relativistic regime for $R\gg R_{\mathrm{N}}(\theta)$. For $R\gg R_{\mathrm{dec}}(\theta)$, the self-similar evolution becomes independent of the value of the initial Lorentz factor $\Gamma_{0}(\theta)$.
Once the Lorentz factor $\Gamma(\theta;R)$ and the velocity $\beta(\theta;R)=\frac{\sqrt{\Gamma^{2}\left(\theta;R\right)-1}}{\Gamma(\theta;R)}$ are known, the corresponding time $t$ in the source frame is given by
@@
This leads to the solution $R(\theta,t)$ and $\Gamma(\theta,t)$ used for the dynamics of each component of the structured jet.
The physical conditions in the shocked medium are easily deduced from the shock jump conditions in the strong shock regime (Blandford & McKee, 1976). In the comoving frame, the mass density equals
@@
We assume here an adiabatic index of $4/3$, valid as long as $\epsilon_{*}\gg c^{2}$. Finally, the timescale of the adiabatic cooling of the shocked region due to the spherical expansion is given in the comoving frame by
@@ 2.3 Accelerated electrons and amplified magnetic field
We consider a shocked region where the physical conditions in the comoving frame are given by the mass density $\rho_{*}$, internal energy per unit mass $\epsilon_{*}$, and dynamical timescale $t^{\prime}_{\mathrm{dyn}}$ (i.e. characteristic timescale of the adiabatic cooling due to the spherical expansion), and we assume that the emission is produced by non-thermal shock-accelerated electrons, that radiate in a local turbulent magnetic field amplified at the shock. In practice, we consider only the forward external shock so that in this study, $\rho_{*}$, $\epsilon_{*}$ and $t^{\prime}_{\mathrm{dyn}}$ are given by Equations (1820).
We use the following standard parametrization of the microphysics at the shock: (i) a fraction $\epsilon_{\mathrm{B}}$ of the internal energy is injected in the magnetic field, i.e.
@@
(ii) a fraction $\epsilon_{\mathrm{e}}$ of the internal energy is injected into non-thermal electrons, that represent a fraction $\zeta$ of all available electrons. Their number density ($\mathrm{cm}^{-3}$) and energy density ($\mathrm{erg\cdot cm^{-3}}$) in the comoving frame are therefore given respectively by
@@
Assuming a power-law distribution at injection, with an index $-p$ and $32<p<$ 3, this leads to the following distribution of accelerated electrons ($\mathrm{cm}^{-3}$):
@@
As discussed later, taking into account a realistic estimate of the maximum Lorentz factor $\gamma_{\mathrm{max}}$ up to which electrons can be accelerated at the shock is important to discuss the GeV-TeV afterglow emission. This was not included in the study by Nakar et al. (2009) on which our calculation of the emission is based. Equations (2526) above are obtained by assuming that the maximum Lorentz factor $\gamma_{\mathrm{max}}$ is much larger than $\gamma_{\mathrm{m}}$. This is fully justified as we have typically $\gamma_{\mathrm{max}}/\gamma_{\mathrm{m}}>10^{6}$ in the best fit models of GW 170817 at the peak (Section 4). The maximum electron Lorentz factor $\gamma_{\mathrm{max}}$ is evaluated by imposing that the acceleration timescale remains always shorter than the radiative and the dynamical timescales, i.e.
@@
where $K_{\rm acc}\geq 1$ is a dimensionless factor. In practice, electrons at $\gamma_{\mathrm{max}}$ are usually fast cooling and the maximum electron Lorentz factor is determined by the radiative timescale, $t^{\prime}_{\mathrm{rad}}(\gamma)$. Its evaluation is non-trivial when Inverse Compton scatterings are taken into account (see § 2.4). The resulting detailed calculation of $\gamma_{\mathrm{max}}$ is explained in Appendix A. In the following, we assume $K_{\mathrm{acc}}=1$. Our value of $\gamma_{\mathrm{max}}$ should therefore be considered as an upper limit for the true value of the maximum electron Lorentz factor.
@@ 2.4 Emissivity in the comoving frame
Our calculation of the emission from non-thermal electrons is mostly taken from Nakar et al. (2009). The population of electrons cools down by adiabatic cooling, synchrotron radiation and via Synchrotron Self-Compton (SSC) scatterings. As we are interested in the VHE afterglow emission, we include in § 2.4.8 the attenuation due to pair production, without including the emission of secondary leptons. As we focus in this paper on the afterglow of GW 170817 where the radio observations do not show any evidence for absorption, we do not include in the present version of the model the effect of synchrotron self-absorption. In this subsection, all quantities are written in the comoving frame of the considered shocked region. We therefore omit the prime to simplify the notations.
2.4.1 Synchrotron and IC radiation for a single electron
@@
with $K_{2}=K_{\nu}\frac{e}{2\pi m_{\mathrm{e}}c}$ and $K_{\nu}$ a dimensionless parameter. The dimensionless parameters $K_{P}$ and $K_{\nu}$ depend on the assumptions made on the pitch angle of electrons relative to the magnetic field lines and different values are found in afterglow models, as listed in Table 1. By default, we use the same values as Gill & Granot (2018), taken from Granot et al. (1999) where they were obtained by fitting the broken power-law approximation of the synchrotron power of a power-law distribution of electrons to the exact calculation.
@@
$K_{P_{\mathrm{max}}}=K_{P}/K_{\nu}$ and $\int_{0}^{\infty}\Phi(x)\mathrm{d}x=1$. In practice we adopt a simplified shape $\Phi(x)=\frac{4}{3}x^{1/3}$ for $x\leq 1$ and 00 otherwise.
Inverse Compton scatterings.
The spectral number density of seed photons is $n_{\nu}^{\mathrm{syn}}$ ($\mathrm{cm^{-3}\cdot Hz^{-1}}$), assumed to result only from the synchrotron radiation of the electrons on which they also scatter (SSC). The IC power of an electron ($\mathrm{erg\cdot s^{-1}}$) is then given by
@@
where $w=\gamma h\nu/m_{\mathrm{e}}c^{2}$; and $f_{\mathrm{KN}}$ and $g_{\mathrm{KN}}$ are the KN corrections to the cross-section and to the mean energy of the scattered photon. In the Thomson regime ($w\ll w_{\mathrm{KN}}$), the cross-section is $\sigma_{\mathrm{T}}$ and the mean energy of scattered photons is $K_{\mathrm{IC}}\,\gamma^{2}h\nu$ for a seed synchrotron photon of frequency $\nu$, so that $f_{\mathrm{KN}}=g_{\mathrm{KN}}=1$ for $w\ll w_{\mathrm{KN}}$. By default, we use the standard values $K_{\mathrm{IC}}=4/3$ and $w_{\mathrm{KN}}=1$, as in Nakar et al. (2009). We note however that Yamasaki & Piran (2022) recently suggested that a better agreement with a detailed numerical spectral calculation of the SSC component was achieved with $w_{\mathrm{KN}}=0.2$, especially at the transition between the Thomson and the KN regime. For simplicity, we do not write explicitly in Equation (33) the dependency on the angle between the upscattered photon and the scattering electron which is discussed in Nakar et al. (2009). We assume an isotropic seed radiation field. We only account for single IC scatterings as the effects of multiple scatterings are expected to be negligible since the afterglow is produced in the optically thin regime. The Compton parameter is defined by
@@
This Compton parameter is constant only if all SSC scatterings occur in the Thomson regime. Here we take into account the KN regime, hence the dependence to the electron Lorentz factor. This affects not only the SSC emission but also the synchrotron component. Following Nakar et al. (2009) and keeping the same notations, we simplify the IC cross-section by assuming that scatterings are entirely suppressed in the KN regime: $f_{\mathrm{KN}}(w)=0$ for $w\geq w_{\mathrm{KN}}$. Then, a synchrotron photon produced by an electron with Lorentz factor $\gamma$ can only be upscattered by electrons with Lorentz factors below a certain limit $\widehat{\gamma}$ to remain in the Thomson regime, i.e.
@@
Note that $\widehat{\gamma}$ is a decreasing function of $\gamma$: if the energy of the synchrotron photon is higher, the maximum energy of any electron on which it could be upscattered in the Thomson regime must be lower. In an equivalent way, electrons with Lorentz factor $\gamma$ will only scatter photons below the energy
@@
Then it is convenient to define $\widetilde{\gamma}$ such that $\nu_{\mathrm{syn}}(\widetilde{\gamma})=\widetilde{\nu}$, i.e.
@@
The definition of $\widetilde{\nu}$ allows to simplify the expression of the SSC power given by Equation (33):
@@
with $u_{\nu}^{\mathrm{syn}}=n_{\nu}^{\mathrm{syn}}\times h\nu$ the energy density of the seed synchrotron photons ($\mathrm{erg\cdot cm^{-3}\cdot Hz^{-1}}$). Then
@@
Assuming a strict suppression in the KN regime leads to a simple expression for the SSC power at frequency $\nu$ ($\mathrm{erg\cdot s^{-1}\cdot Hz^{-1}}$) for a single electron of Lorentz factor $\gamma$:

### v2

@@ 2.2 Dynamics of the deceleration
The structured outflow decelerates in an external medium, with an assumed density profile as a function of radius $R$,
@@
with $s=0$ for a constant-density external medium as considered in the following to model the afterglow of GW 170817. In this case, we write $A=n_{\rm ext}m_{\rm p}$, with $n_{\rm ext}$ the external medium density. We focus on the dynamics of the shocked external medium at the forward shock. It is computed assuming that (i) the dynamics of each ring of material at angle $\theta$ is independent of other angles, and that (ii) the lateral expansion of the outflow is negligible. These two assumptions are questionable at late times, close to the transition to the Newtonian regime (Rhoads 1997).
The Lorentz factor $\Gamma(\theta;R)$ of the shocked material at angle $\theta$ and radius $R$ is computed in a simplified way, using energy conservation (see e.g. Panaitescu & Kumar 2000; Gill & Granot 2018),
@@
with $M_{\mathrm{ext}}(R)=\int_{0}^{R}\rho_{\mathrm{ext}}(R)\,R^{2}\mathrm{d}R$ the swept-up mass per unit solid angle, and $M_{\mathrm{ej}}(\theta)=\epsilon_{0}(\theta)/\Gamma_{0}(\theta)c^{2}$ the ejected mass per unit solid angle. The deceleration radius of the material at angle $\theta$ is defined by
@@
For a power-law structure as defined by Eqs. (5) and (8), the deceleration radius scales as $R_{\mathrm{dec}}(\theta)\propto\theta^{\frac{2b-a}{3-s}}$. Best-fit models of the afterglow of GW 170817 usually have $2b>a$ so that the deceleration radius is larger at high latitude (see Sect. 4.2). More generally, Beniamini et al. (2020b) derived the conditions for the observed afterglow light curve to have a single peak, as in the case of GW 170817. For an off-axis observation with a large viewing angle, these conditions are $\Gamma_{0}^{\mathrm{c}}\theta_{\mathrm{c}}>1$; $2b>a/(4-s)$ (i.e. $8b>a$ for a uniform external medium), $b>-\ln{\Gamma_{0}^{\mathrm{c}}}/\ln{\theta_{\mathrm{c}}}$, and $\theta_{\mathrm{v}}>\theta_{\mathrm{c}}\left(\Gamma_{0}^{\mathrm{c}}\theta_{\mathrm{c}}\right)^{\frac{1}{b-1}}$.
This description allows us to characterise the early-time dynamics in the coasting phase $R\ll R_{\mathrm{dec}}(\theta)$ where the Lorentz factor remains constant. It continuously branches out to the relativistic self-similar evolution (Blandford & McKee 1976) for $R_{\mathrm{dec}}(\theta)\ll R\ll R_{\mathrm{N}}(\theta)=\Gamma_{0}^{\frac{2}{3-s}}(\theta)R_{\mathrm{dec}}(\theta)$, and finally, to the Sedov-Taylor phase (Sedov 1946; Taylor 1950) in the non-relativistic regime for $R\gg R_{\mathrm{N}}(\theta)$. For $R\gg R_{\mathrm{dec}}(\theta)$, the self-similar evolution becomes independent of the value of the initial Lorentz factor $\Gamma_{0}(\theta)$.
When the Lorentz factor $\Gamma(\theta;R)$ and the velocity $\beta(\theta;R)=\frac{\sqrt{\Gamma^{2}\left(\theta;R\right)-1}}{\Gamma(\theta;R)}$ are known, the corresponding time $t$ in the source frame is given by
@@
This leads to the solution $R(\theta,t)$ and $\Gamma(\theta,t)$ that we use for the dynamics of each component of the structured jet.
The physical conditions in the shocked medium are easily deduced from the shock-jump conditions in the strong-shock regime (Blandford & McKee 1976). In the comoving frame, the mass density equals
@@
We assumed an adiabatic index of $4/3$, which is valid as long as $\epsilon_{*}\gg c^{2}$. Finally, the timescale of the adiabatic cooling of the shocked region due to the spherical expansion is given in the comoving frame by
@@ 2.3 Accelerated electrons and amplified magnetic field
We consider a shocked region where the physical conditions in the comoving frame are given by the mass density $\rho_{*}$, the internal energy per unit mass $\epsilon_{*}$, and the dynamical timescale $t^{\prime}_{\mathrm{dyn}}$ (i.e. the characteristic timescale of the adiabatic cooling due to the spherical expansion), and we assume that the emission is produced by non-thermal shock-accelerated electrons that radiate in a local turbulent magnetic field amplified at the shock. In practice, we only consider the forward external shock, so that in this study, $\rho_{*}$, $\epsilon_{*}$, and $t^{\prime}_{\mathrm{dyn}}$ are given by Eqs. (18)–(20).
We use the following standard parametrisation of the microphysics at the shock: (i) a fraction $\epsilon_{\mathrm{B}}$ of the internal energy is injected in the magnetic field,
@@
(ii) a fraction $\epsilon_{\mathrm{e}}$ of the internal energy is injected into non-thermal electrons that represent a fraction $\zeta$ of all available electrons. Their number density ($\mathrm{cm}^{-3}$) and energy density ($\mathrm{erg\cdot cm^{-3}}$) in the comoving frame are therefore given by
@@
We assume a power-law distribution at injection with an index $-p$ and $32<p<$ 3. This leads to the following distribution of accelerated electrons ($\mathrm{cm}^{-3}$):
@@
As discussed below, it is important to take a realistic estimate of the maximum Lorentz factor $\gamma_{\mathrm{max}}$ up to which electrons can be accelerated at the shock into account for a discussion of the GeV-TeV afterglow emission. This was not included in the study by Nakar et al. (2009), on which we base our calculation of the emission. Eqs. (25) and (26) above were obtained by assuming that the maximum Lorentz factor $\gamma_{\mathrm{max}}$ is much higher than $\gamma_{\mathrm{m}}$. This is fully justified as we have typically $\gamma_{\mathrm{max}}/\gamma_{\mathrm{m}}>10^{6}$ in the best-fit models of GW 170817 at the peak (Sect. 4). The maximum electron Lorentz factor $\gamma_{\mathrm{max}}$ is evaluated by imposing that the acceleration timescale always remains shorter than the radiative and dynamical timescales,
@@
where $K_{\rm acc}\geq 1$ is a dimensionless factor. In practice, electrons at $\gamma_{\mathrm{max}}$ usually cool fast, and the maximum electron Lorentz factor is determined by the radiative timescale, $t^{\prime}_{\mathrm{rad}}(\gamma)$. Its evaluation is non-trivial when IC scattering is taken into account (see Sect. 2.4). The resulting detailed calculation of $\gamma_{\mathrm{max}}$ is explained in Appendix A. In the following, we assume $K_{\mathrm{acc}}=1$. Our value of $\gamma_{\mathrm{max}}$ should therefore be considered as an upper limit for the true value of the maximum electron Lorentz factor.
@@ 2.4 Emissivity in the comoving frame
Our calculation of the emission from non-thermal electrons is mostly taken from Nakar et al. (2009). The population of electrons cools down by adiabatic cooling, synchrotron radiation, and via SSC scattering. As we are interested in the VHE afterglow emission, we include the attenuation due to pair production in Sect. 2.4.8, but do not include the emission of secondary leptons. We focus on the afterglow of GW 170817, for which the radio observations did not show any evidence for absorption, and we therefore do not include the effect of synchrotron self-absorption in the current version of the model. In this subsection, all quantities are written in the comoving frame of the considered shocked region. We therefore omit the prime to simplify the notations.
2.4.1 Synchrotron and inverse-Compton radiation for a single electron
@@
with $K_{2}=K_{\nu}\frac{e}{2\pi m_{\mathrm{e}}c}$ and $K_{\nu}$ a dimensionless parameter. The dimensionless parameters $K_{P}$ and $K_{\nu}$ depend on the assumptions made on the pitch angle of electrons relative to the magnetic field lines and different values are found in afterglow models, as listed in Table 1. By default, we use the same values as Gill & Granot (2018), taken from Granot et al. (1999), who obtained them by fitting the broken power-law approximation of the synchrotron power of a power-law distribution of electrons to the exact calculation.
@@
$K_{P_{\mathrm{max}}}=K_{P}/K_{\nu}$ and $\int_{0}^{\infty}\Phi(x)\mathrm{d}x=1$. In practice, we adopt a simplified shape $\Phi(x)=\frac{4}{3}x^{1/3}$ for $x\leq 1$ and 00 otherwise.
Inverse-Compton scattering.
The spectral number density of seed photons is $n_{\nu}^{\mathrm{syn}}$ ($\mathrm{cm^{-3}\cdot Hz^{-1}}$), which is assumed to result only from the synchrotron radiation of the electrons on which they also scatter (SSC). The IC power of an electron ($\mathrm{erg\cdot s^{-1}}$) is then given by
@@
where $w=\gamma h\nu/m_{\mathrm{e}}c^{2}$; and $f_{\mathrm{KN}}$ and $g_{\mathrm{KN}}$ are the KN corrections to the cross-section and to the mean energy of the scattered photon. In the Thomson regime ($w\ll w_{\mathrm{KN}}$), the cross-section is $\sigma_{\mathrm{T}}$ and the mean energy of scattered photons is $K_{\mathrm{IC}}\,\gamma^{2}h\nu$ for a seed synchrotron photon with a frequency $\nu$, so that $f_{\mathrm{KN}}=g_{\mathrm{KN}}=1$ for $w\ll w_{\mathrm{KN}}$. By default, we use the standard values $K_{\mathrm{IC}}=4/3$ and $w_{\mathrm{KN}}=1$, as in Nakar et al. (2009). We note, however, that Yamasaki & Piran (2022) recently suggested that a better agreement with a detailed numerical spectral calculation of the SSC component was achieved with $w_{\mathrm{KN}}=0.2$, especially at the transition between the Thomson and the KN regime. For simplicity, we do not write in Eq. (33) the dependence on the angle between the upscattered photon and the scattering electron explicitly. This was discussed in Nakar et al. (2009). We assume an isotropic seed radiation field. We only account for single IC scattering as the effects of multiple scatterings are expected to be negligible because the afterglow is produced in the optically thin regime. The Compton parameter is defined by
@@
This Compton parameter is constant only when SSC scattering occurs in the Thomson regime. We take the KN regime into account, which explains the dependence on the electron Lorentz factor. This affects not only the SSC emission, but also the synchrotron component. Following Nakar et al. (2009) and keeping the same notations, we simplify the IC cross-section by assuming that scattering is entirely suppressed in the KN regime: $f_{\mathrm{KN}}(w)=0$ for $w\geq w_{\mathrm{KN}}$. Then, a synchrotron photon produced by an electron with Lorentz factor $\gamma$ can only be upscattered by electrons with Lorentz factors below a certain limit $\widehat{\gamma}$ to remain in the Thomson regime,
@@
We note that $\widehat{\gamma}$ is a decreasing function of $\gamma$: If the energy of the synchrotron photon is higher, the maximum energy of any electron on which it could be upscattered in the Thomson regime must be lower. In an equivalent way, electrons with Lorentz factor $\gamma$ only scatter photons below the energy
@@
Then, it is convenient to define $\widetilde{\gamma}$ such that $\nu_{\mathrm{syn}}(\widetilde{\gamma})=\widetilde{\nu}$,
@@
The definition of $\widetilde{\nu}$ allows us to simplify the expression of the SSC power given by Eq. (33):
@@
with $u_{\nu}^{\mathrm{syn}}=n_{\nu}^{\mathrm{syn}}\times h\nu$ the energy density of the seed synchrotron photons ($\mathrm{erg\cdot cm^{-3}\cdot Hz^{-1}}$). Then,
@@
Assuming a strict suppression in the KN regime leads to a simple expression for the SSC power at frequency $\nu$ ($\mathrm{erg\cdot s^{-1}\cdot Hz^{-1}}$) for a single electron with a Lorentz factor $\gamma$,